OpenStudy (anonymous):

what is u=g(x) of the the function e^5(sqrtx)? I though its Sqrtx but its not

OpenStudy (anonymous):

\[e^5\sqrt{x}\]

OpenStudy (anonymous):

What are you trying to do with it malik? :P

OpenStudy (anonymous):

never mind its \[5\sqrt{x}\]

OpenStudy (anonymous):

i was trying to figure out the u form

OpenStudy (anonymous):

but wait can you help how to do the dy/dx? please

OpenStudy (anonymous):

Of course :) The 5 is a constant so you can pull that out of the derivative. 5(d/dx)x^(1/2) (rewriting for convenience and demonstrative purposes) Its just a power rule then: 5(1/2)x^(-1/2) Or: \[\frac{5}{2}x^{\frac{-1}{2}}=\frac{5}{2\sqrt{x}}\]

OpenStudy (anonymous):

i got the same!! im so happy i finally understand..thanks a lot!!! your a life saver!!!

OpenStudy (anonymous):

Haha, no problem malik :P

OpenStudy (anonymous):

wait but what happens to the e?

OpenStudy (anonymous):

You said "never mind its 5sqrt(x)". If you rewrite it with an e we can do that one too.

OpenStudy (anonymous):

ok hold on....\[e^5\sqrt{x}*5\sqrt{x}\]

OpenStudy (anonymous):

so we already know the second equation but how to compute the first derivatives

OpenStudy (anonymous):

Okay: If they are multiplied then you have: \[5e^5*\sqrt{x}*\sqrt{x}=5e^5x\] \[5e^5 \frac{d}{dx}(x)=5e^5\]

OpenStudy (anonymous):

ok i think i ask the wrong question haha sorry...i have to dy/dx the \[e^5\sqrt{x}\]

OpenStudy (anonymous):

and it told me to find the u=g(x) and g=f(u)..which we've found...